# Building spatial world models from sparse transitional episodic memories

### v1

@@ A.7 Baseline Navigational Agent
Appendix B Additional Figures
@@
Figure A1: ISOMAP projection of various ESWM-trained models’ latent space. The last two environments are out-of-distribution — ESWM has never seen multiple walls in the same room nor a cluster of obstacles that does not form a long, straight wall during training. For Transformer-2L, LSTM-4L, and Random network (LSTM-4L initialized with random weights), the activations for end-state prediction are used while for Mamba, the activations for action prediction are used. EPN (Section D) is an episodic-memory-based navigational and exploration agent trained on the same environment as ESWM.
Figure A2: Same ISOMAP projections as Fig.A1 but in a different angle.
@@
Figure A3: ESWM’s latent spatial map persists for all prediction tasks and partial observations. a) ESWM’s latent spatial map for start-state, action, and end-state predictions. Mamba is excluded as the spatial map is only found for the action prediction task. For Transformer-2L, the spatial map is found in the first layer for action prediction, and the second layer for start-state and end-state prediction. b) Spatial map when the memory bank observes an irregular subsection of the room. Transformer-2L’s first layer’s activation for action prediction is used.
@@
Figure A4: ESWM’s latent spatial map adjusts to structural changes. ISOMAP projections of the 14-layer transformer model’s seventh layer’s activations. a) ESWM’s latent space when the input memory bank $M$ consists of two disjoint memory clusters. b) ESWM’s latent space when both memory clusters observe a wall to their left; despite remaining separate in the memory bank, ESWM inferred them to be connected as reflected in its latent space. c) ESWM’s latent space when both memory clusters observe a single obstacle to their left. ESWM, again, inferred the clusters are connected but with sparser observations of obstacles than in b). 900 activations and 45 neighbours are used to fit the ISOMAP.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure A5: Representation becomes more continuous with depth. Left to right are 30-neighbours ISOMAP projections from layer 1, 6, 9, and 14 of Transformer-14L (2000 activations). The model was trained without random masking of subregions. Representation in the latent space becomes progressively more continuous.
Figure A6: Exploration, navigation, and adaptability for ESWM and EPN. a) Number of unique states visited by different exploration strategies in a Random Wall environment with 37 locations over 25 time steps (n=1000). On average, ESWM explores a comparable number of states to an oracle explorer with access to obstacle locations. b, c, d) Performance on Random Wall with 19 locations of individual EPN’s random seeds, and EPN with memory bank size ablated from 200 to 50 (small memory). b) Number of unique states visited over the initial 15 time steps in a novel environment. c) Navigation success rate and optimality when tasked to navigate to a goal across various path lengths, n = 2400 paths. d) Adaptation to structural changes in the environment. Success rate of navigating to a goal location when obstacles contradicting past experiences are inserted
@@
Figure A7: The latent space of ESWM can be harnessed as a heuristic for more efficient planning. All the results are on Random Wall environment with 37 locations. a) (Left) Average number of nodes expanded for different heuristics over 1000 paths. (Middle) Success rates of greedy navigation using different heuristics over 2400 paths. (Right) Success rates for greedy navigation on paths requiring detours around obstacles over 800 paths. b) (Left) Correlation between geodesic distances on $G_{\text{latent}}$ (see 4.6.2) and corresponding distances in the environment, using n=1500 pairs of states $(s_{\text{start}},s_{\text{goal}})$ over 75 environments. The linear fit shows a strong positive correlation with $R^{2}=0.89$. (Mid-Left) sample grid environment. (Mid-Right) $G_{\text{latent}}$ for the same environment. (Right) Agent navigates greedily on $h_{\text{latent}}$ on a path that requires detours around obstacles. c) Comparison of navigation success rate and path optimality between Dijkstra’s and A* over different path lengths (n=1500). Although not guaranteed to be admissible, $h_{\text{latent}}$ empirically preserves path optimality.
Figure A8: Comparison of end-state prediction accuracy between EMWM Coda-Forno et al. [2022a] and Transformer-2L ESWM in RandomWall. EMWM predicts using a weighted average over the top-K matches (K=5), but we found K=1 performs best, likely due to the sparsity of our memory banks (i.e, transitions are unique, so averaging more than one transition offers no benefit). To adapt the author’s implementation to our task, we embed each token (transitional memory tokens or the end-state masked query) by averaging the embedding of each item ($s_{t},a_{t},s_{t+1}$), each produced using a distinct MLP layer – identical to ESWM. All tokens are sequentially fed into a 1-layer GRU Cho et al. [2014] to obtain the sequence of hidden states, which become the keys with the corresponding end-state embeddings as values for the memory table. The network is optimized to predict the $s_{t+1}$ for each token.
Figure A9: ESWM in scaled-up MiniGrid environments. (Top left) An example of a procedurally-generated 9$\times$9 environment with random coloring patterns in which ESWM is trained. Episodic memory is considered as a transition between 5x5 egocentric pixel observations via egocentric actions (turn left, turn right, and move forward). Memory bank is a set of one-step transitions that minimally observe the entire room, with some degree of overlap between transitions to enable memory integration. (Top right) Accuracies for start-state, action, and end-state predictions. (Bottom) ISOMAP analysis on ESWM’s latent space shows a map-like representation.
Appendix C Algorithms
@@
[…]
3:     Initialize priority queue $OpenSet$
4:     Initialize dictionary $Cost$
5:     Initialize dictionary $Parent$
@@
7:     Push $(s_{\text{start}},\,cost=0,\,depth=0)$ into $OpenSet$
@@
11:              return ReconstructPath($Parent$, $u$)
@@
21:                  Push $(v,\,c^{\prime},\,d^{\prime})$ into $OpenSet$
@@
[…]
7:              Enqueue $a$ into $q$ with $\text{cost}=1-\beta\textsc{Entropy}(\textit{probs}))$triangleright $\beta$ = 0.2 in our experiments
@@
9:         nodesToSAs[$activation$] $\leftarrow(s,a)$
10:         activations.append($activation$)
@@
16:              $h_{\text{latent}}$[$sa_{i},sa_{j}$] $\leftarrow$ ShortestPathLength($G_{\text{latent}}$, $n_{i},n_{j}$)
@@
Appendix D Episodic Planning Networks (EPN)
The Episodic Planning Networks (EPN) model Ritter et al. [2020] is specifically designed and trained to rapidly learn to navigate in novel environments. Comparatively, ESWM is a general-purpose model not explicitly trained for navigation. By comparing with EPN, we investigate whether ESWM’s ability to construct internal models of spatial structures enables navigational capabilities. Similarly to ESWM, EPN is designed to leverage episodic memory to facilitate rapid adaptation in unfamiliar environments, and both models are episodic memory-driven transformer-based models.
Hyperparameters – We use the same hyperparameters as those reported by Ritter et al. [2020].
Training – Both models are trained on the same environment structure variations and amount of data, approximately $91.14e$ 9 steps, which is a restrictive data regime for EPN. EPN is trained to navigate through reinforcement learning, using IMPALA [Espeholt et al., 2018]. ESWM is trained to build internal model of environments from sparse disjointed episodic experiences through self-supervision.
Memory bank – EPN has a memory bank containing 200 transitions, whereas ESWM only stores the minimal spanning tree, with 19 transitions for smaller environments and 36 for larger ones. When evaluating EPN ability to navigate, we follows the methodology from Ritter et al. [2020]. EPN explores to populate its memory bank for the first 2/3 of episodes , and the performance is only measured on the last third.
Appendix E Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, particularly in autonomous exploration, navigation, and decision-making in real-world environments. While our work primarily focuses on theoretical and methodological advancements, potential societal consequences include enhanced autonomous navigation technologies, more efficient exploration strategies in robotics, and improved AI models inspired by human cognitive abilities. We also acknowledge the potential misuse of the technology in collecting sensitive data and the ethical considerations regarding the deployment of autonomous agents in critical decision-making scenario, emphasizing the importance of responsible and safe integration of such models in real-world applications.

### v2

@@ C.10 Baseline Navigational Agent
Appendix D Additional Analysis
D.1 Comparing Models in Open Arena
Different models exhibit drastically different performance in Open Arena. This section aims to provide insights into the factors that contribute to each model’s performance.
The ESWM-T models perform noticeably well in the Open Arena. We investigate which hyperparameter affect its performance the most. In Fig.2, we fix the number of heads and observe a positive correlation between model depth and performance. In Table. 11, we fix the depth and vary the number of heads. We observe that while increasing depth leads to a monotonic increase in performance, increasing the number of heads initially improves performance, but exceeding 8 heads leads to a quick drop in performance. In addition, the model depth impacts all three prediction tasks, while the number of heads selectively impacts state prediction tasks more than the action prediction task.
As shown in Fig.8, ESWM-MAMBA and ESWM-LSTM both overfit on the training states, explaining their poor performance in completely unseen test states. However, we note an additional observation: LSTM struggles to perform even on the training set in Open Arena, unlike in Random Wall where it succeeds on both train and evaluation splits. We suspect another task difference between Open Arena and Random Wall contributes to this discrepancy: states in Open Arena consist of a 6-bit value, each represented by a 128-dimensional vector while states in Random Wall consist of a single continuous vector. The latter may allow a more compact latent representation that better fits within LSTM’s limited memory capacity.
TEM-T models, which also incorporate transformer into its architecture, underperform compared with ESWM-T models. TEM-T is architecturally distinct (See Section 2) and is trained on a different distribution (minimal spanning trajectories rather than minimal spanning trees) compared to ESWM-T. To determine which of these two factors contributes to the performance gap, we train ESWM-T-10L on TEM-T’s training distribution and find that it can reach a similar performance compared with training on ESWM’s, predicting unseen $s_{s}$, $a$, $s_{e}$ with accuracies 99%, 99%, 99%. The similar is true in Random Wall, where ESWM-T-6L reaches $s_{s}$, $a$, $s_{e}$ accuracies of 81%, 89%, 78% when trained with TEM-T’s training distribution. These results indicate that TEM-T’s architecture most likely explains its performance gap with ESWM-T.
D.2 Comparing ESWM Trained With or Without the Minimality Constraint on Memory Banks
We show that ESWM-T performs well when trained with minimal memory banks (Fig. 2) and non-minimal memory banks (Fig. 9). However, ESWM-T trained on non-minimal memory banks, where more environmental transitions are observed, underpeforms the one trained on minimal memory bank, where less environmental transitions are observed.
To investigate why the ESWM-T trained on non-minimal memory banks underperforms while having access to more experience in an environment, we break down the model’s performance across memory bank sizes and observe a substantial gap between denser and sparser memory banks, even though they are equally likely to be in the training data (Fig. 9). This suggests that, when trained on a mixture of densities, the model focuses more on solving the easier tasks where it is given an overcomplete memory bank, and only partially succeeds in learning the more challenging tasks when minimal information is given. This leads to weaker spatial reasoning skills and poorer performance when the observation pattern changes.
In contrast, the ESWM-T that is trained under minimality constraints is forced to solve the harder spatial reasoning problem consistently. As shown in the (Fig. 4c) this produces representations that transfer robustly to denser, Out-of-Distribution (OOD) memory banks. We also note that within the training distribution (ID), increasing memory bank size corresponds to an expansion of the observed area, which requires the model to coherently build a larger map which is more challenging and likely explains the performance drop.
Collectively, these findings suggest that applying sparsity constraints during training leads to better test-time flexibility in terms of information availability.
D.3 Handling Disconnected Memory Clusters
During Random Wall training, memory banks can contain disconnected memory clusters, with no paths between observations in distinct clusters. This leads to uncertainty in spatial relationships as path integration provides no avail. In rare cases, the model is asked to solve queries that bridge disconnected clusters. This raises two questions: (1) Does the model infer the relative positions of disconnected memory clusters (e.g., which part of the room each cluster maps onto)? (2) Which cues enable these inferences?
D.3.1 Inference from Environmental Structure
We observe that when memory clusters are disconnected, the model leverages the underlying structure of the room to infer their relative positions. For example, the model aligns memory clusters with the room’s boundary when their shape matches. This allows the model to piece together disconnected memory clusters, like solving a puzzle, by identifying the configuration that satisfies the invariant structural constraints. Evidence for this behaviour is twofold. First, in the model’s latent space, the disconnected clusters are stitched together, rather than overlapping or disconnecting, to form a smooth surface reflecting the room’s structure (Fig. 3c). Second, the model’s end state prediction distributions, starting from the boundary states of one cluster, are biased towards the boundary states of the inferred neighbouring cluster, instead of exclusively outputting the “I don’t know” option.
D.3.2 Generalization from Observations of Obstacles
In cases where disconnected memory clusters’ shapes do not provide clear cues about their relative locations, the model relies on observations of obstacles to deduce their positions. For instance, given two disconnected clusters of memories, the model initially maps them to different locations in its latent space and outputs predictions exclusively reflecting uncertainty (“I don’t know”). However, if both clusters observe obstacles, even though the clusters remain disconnected, the model applies its knowledge of common wall shapes (e.g, walls are usually straight) to make inferences about the relative locations of the clusters. In the latent space, the two previously disconnected clusters are connected and aligned along the inferred wall shape to form a continuous representation (Fig. 3d, 13). The model’s predictions are also biased towards the inferred neighbours. Moreover, a single observation of obstacles from both clusters is sufficient for the model to infer the shape of the wall, and therefore infer the relative positions of the clusters (Fig. 13). These results demonstrate that ESWM can leverage the invariant structure of the room and combine observations of obstacles with prior knowledge of wall shapes to construct a coherent spatial representation.
D.4 Planning by Guided Imagination
While the strategy introduced in the main paper enables planning, it uniformly searches through the space, unbiasedly imagining the consequences of all actions, even if some move further from the goal. In contrast, animals with internal maps encoding geometric constraints plan biasedly, prioritizing actions that shorten the geodesic distance (i.e, the shortest path length) to the goal. We show that the spatial maps in ESWM’s latent space, which accurately encode obstacles and geodesic distance between states (Fig. 16b), can be used for guided imagination towards the goal.
The previous strategy is akin to Dijkstra’s algorithm, or A* with a null heuristic function (Hart et al., 1968; Ferguson et al., 2005). The key to efficient planning, which prunes undesirable actions early, then, is a heuristic function containing rich information on the geodesic distances between states in the environment. Inspired by ISOMAP Tenenbaum et al. (2000), we construct such a heuristic function $h_{\text{latent}}$ as follows (See Alg.4 for pseudocode): 1) Query the model to solve the end state prediction task for all possible $(s,a)$ pairs, collecting activations. 2) Construct a graph $G_{\text{latent}}$ where nodes are activations and edges connect nodes within radius $R_{\text{latent}}$ (Appendix 4). Edges are weighted by the cosine distance between their two endpoints (Fig. 16b). Intuitively, each node on $G_{\text{latent}}$ is the activation of predicting the outcome of action $a$ starting at state $s$. 3) Compute pairwise shortest paths on this graph to create a geodesic distance table $h_{\text{latent}}$.
$h_{\text{latent}}$ captures ESWM’s latent spatial map into a look-up table containing information on which action takes you closer to the goal state at every step of imagination. A* guided imagination with $h_{\text{latent}}$, improves planning efficiency by 57% (Fig. 16a), as measured by the number of states visited, over Dijkstra’s while maintaining path optimality (Fig. 16c). Furthermore, greedy policy based on $h_{\text{latent}}$ achieves 70% success in navigating around obstacles (Fig. 16a). These results suggest that structured latent spaces in world models can serve as efficient heuristics for planning.
Appendix E Additional Tables
2 Heads 4 Heads 8 Heads 16 Heads
Predict Source 40.4 78.9 83.3 42.0
Predict Action 84.6 91.8 92.6 91.8
Predict End 37.6 78.1 79.6 40.4
Table 1: ESWM-T number of heads ablation
. We report the accuracies for the start state, action, and end state prediction task in Open Arena with ESWM-T-6L.
Predict Source Predict Action Predict End
Seen 0.9050 0.8700 1.0000
Unseen 0.8750 0.8880 0.9030
Unsolvable 0.8270 0.9050 0.8180
Table 2: ESWM-T-4L performance in Random Wall
. Accuracies for source, action, and end prediction.
Predict Source Predict Action Predict End
Seen 0.8780 0.8690 1.0000
Unseen 0.6660 0.7530 0.7030
Unsolvable 0.7620 0.8830 0.7560
Table 3: ESWM-LSTM-4L performance in Random Wall
. Accuracies for source, action, and end prediction.
Predict Source Predict Action Predict End
Seen 0.9000 0.8810 1.0000
Unseen 0.7710 0.8110 0.8230
Unsolvable 0.8110 0.8990 0.8010
Table 4: ESWM-Mamba-5L performance in Random Wall
. Accuracies for source, action, and end prediction.
Appendix F Additional Figures
Figure 7: ESWM’s evaluation accuracy in grid environments with duplicated states. A trajectory is prepended to the query transition to serve as context for disambiguating two locations with the same state. The context trajectory begins at a random state and ends at $s_{s}$ of the query transition. The query transition’s $a$ and $s_{e}$ are masked with equal chances. a) Accuracy in Open Arena for ESWM-T-8L and b) Random Wall for ESWM-T-4L.
Figure 8: Training and evaluation loss for various ESWM models in Open Arena.
Figure 9: Training with non-minimal memory banks.
a) Evaluation accuracy on unseen queries for various models trained with non-minimal memory banks. ESWM receives memories that are disjoint, spanning, but not minimal (constructed by adding extra episodic memories to the minimal memory banks) and is tasked to predict an unseen query. TEM receives a minimal spanning trajectory padded with random walks as input. b) Prediction accuracy over memory bank sizes for ESWM-T-12L from a). The shaded area emphasizes that this model is trained on non-minimal memory banks. ID means In Distribution.
@@ Appendix F Additional Figures
Figure 10: ISOMAP projection of various ESWM-trained models’ latent space. The last two environments are out-of-distribution — ESWM has never seen multiple walls in the same room nor a cluster of obstacles that does not form a long, straight wall during training. For Transformer-2L, LSTM-4L, and Random network (LSTM-4L initialized with random weights), the activations for end-state prediction are used while for Mamba, the activations for action prediction are used. EPN (Section H) is an episodic-memory-based navigational and exploration agent trained on the same environment as ESWM.
Figure 11: Same ISOMAP projections as Fig. 10 but in a different angle.
@@
Figure 12: ESWM’s latent spatial map persists for all prediction tasks and partial observations. a) ESWM’s latent spatial map for $s_{s}$, $a$, and $s_{e}$ predictions. Mamba is excluded as the spatial map is only found for the action prediction task. For Transformer-2L, the spatial map is found in the first layer for $a$ prediction, and the second layer for $s_{s}$ and $s_{e}$ prediction. b) Spatial map when the memory bank observes an irregular subsection of the room. Transformer-2L’s first layer’s activation for action prediction is used.
@@
Figure 13: ESWM’s latent spatial map adjusts to structural changes. ISOMAP projections of the 14-layer transformer model’s seventh layer’s activations. a) ESWM’s latent space when the input memory bank $M$ consists of two disjoint memory clusters. b) ESWM’s latent space when both memory clusters observe a wall to their left; despite remaining separate in the memory bank, ESWM inferred them to be connected as reflected in its latent space. c) ESWM’s latent space when both memory clusters observe a single obstacle to their left. ESWM, again, inferred the clusters are connected but with sparser observations of obstacles than in b). 900 activations and 45 neighbours are used to fit the ISOMAP.
Figure 14: Representation becomes more continuous with depth. Left to right are 30-neighbours ISOMAP projections from layer 1, 6, 9, and 14 of Transformer-14L (2000 activations). The model was trained without random masking of subregions. Representation in the latent space becomes progressively more continuous.
Figure 15: Exploration, navigation, and adaptability for ESWM and EPN. a) Number of unique states visited by different exploration strategies in a Random Wall environment with 37 locations over 25 time steps (n=1000). On average, ESWM explores a comparable number of states to an oracle explorer with access to obstacle locations. b, c, d) Same as Fig. 5 except the performance of ESWM and EPN is averaged over 4 random seeds. Error bars show standard deviation across seeds. A Welch’s two-sample t-test was used to compare ESWM and EPN ($\alpha$ = 0.05). A dashed line with “*” indicates a statistically significant difference (p<p<0.05)
@@
Figure 16: The latent space of ESWM can be harnessed as a heuristic for more efficient planning. All the results are on Random Wall environment with 37 locations. a) (Left) Average number of nodes expanded for different heuristics over 1000 paths. (Middle) Success rates of greedy navigation using different heuristics over 2400 paths. (Right) Success rates for greedy navigation on paths requiring detours around obstacles over 800 paths. b) (Left) Correlation between geodesic distances on $G_{\text{latent}}$ (see D.4) and corresponding distances in the environment, using n=1500 pairs of states $(s_{\text{start}},s_{\text{goal}})$ over 75 environments. The linear fit shows a strong positive correlation with $R^{2}=0.89$. (Mid-Left) sample grid environment. (Mid-Right) $G_{\text{latent}}$ for the same environment. (Right) Agent navigates greedily on $h_{\text{latent}}$ on a path that requires detours around obstacles. c) Comparison of navigation success rate and path optimality between Dijkstra’s and A* over different path lengths (n=1500). Although not guaranteed to be admissible, $h_{\text{latent}}$ empirically preserves path optimality.
Figure 17: Comparison of end-state prediction accuracy between EMWM Coda-Forno et al. (2022a) and Transformer-2L ESWM in RandomWall. EMWM predicts using a weighted average over the top-K matches (K=5), but we found K=1 performs best, likely due to the sparsity of our memory banks (i.e, transitions are unique, so averaging more than one transition offers no benefit). To adapt the author’s implementation to our task, we embed each token (transitional memory tokens or the end-state masked query) by averaging the embedding of each item ($s_{t},a_{t},s_{t+1}$), each produced using a distinct MLP layer – identical to ESWM. All tokens are sequentially fed into a 1-layer GRU Cho et al. (2014) to obtain the sequence of hidden states, which become the keys with the corresponding end-state embeddings as values for the memory table. The network is optimized to predict the $s_{t+1}$ for each token.
Appendix G Algorithms
@@
2:  constACTION_COSTACTION\_COST
3:  Initialize priority queue OpenSetOpenSet
4:  Initialize dictionary CostCost
5:  Initialize dictionary ParentParent
@@
7:  Push $(s_{\text{start}},\,cost=0,\,depth=0)$ into OpenSetOpenSet
@@
11:     return ReconstructPath(ParentParent, $u$)
@@
21:      Push $(v,\,c^{\prime},\,d^{\prime})$ into OpenSetOpenSet
@@
5:   $\textit{probs}\leftarrow$ softmax(logits)
6:   if $s^{\prime}=\text{'I don’t know'}$ then
7:     Enqueue $a$ into $q$ with $\text{cost}=1-\beta\textsc{Entropy}(\textit{probs}))$ $\triangleright$ $\beta$ = 0.2 in our experiments
@@
9:   nodesToSAs[activationactivation] $\leftarrow(s,a)$
10:   activations.append(activationactivation)
@@
16:     $h_{\text{latent}}$$sa_{i},sa_{j}$] $\leftarrow$ ShortestPathLength($G_{\text{latent}}$, $n_{i},n_{j}$)
@@
Appendix H Episodic Planning Networks (EPN)
The Episodic Planning Networks (EPN) model Ritter et al. (2020) is specifically designed and trained to rapidly learn to navigate in novel environments. Comparatively, ESWM is a general-purpose model not explicitly trained for navigation. By comparing with EPN, we investigate whether ESWM’s ability to construct internal models of spatial structures enables navigational capabilities. Similarly to ESWM, EPN is designed to leverage episodic memory to facilitate rapid adaptation in unfamiliar environments, and both models are episodic memory-driven transformer-based models.
Hyperparameters – We use the same hyperparameters as those reported by Ritter et al. (2020).
Training – Both models are trained on the same environment structure variations and amount of data, approximately $1.14e9$ steps, which is a restrictive data regime for EPN. EPN is trained to navigate through reinforcement learning, using IMPALA (Espeholt et al., 2018). ESWM is trained to build internal model of environments from sparse disjointed episodic experiences through self-supervision.
Memory bank – EPN has a memory bank containing 200 transitions, whereas ESWM only stores the minimal spanning tree, with 19 transitions for smaller environments and 36 for larger ones. When evaluating EPN ability to navigate, we follow the methodology from Ritter et al. (2020). EPN explores to populate its memory bank for the first 2/3 of episodes , and the performance is only measured on the last third.
Appendix I Impact Statements
This paper aims to advance the foundations of machine learning by developing a model for structured memory and spatial understanding. Although our primary contributions are theoretical and methodological, the ideas explored here connect directly to real-world applications in autonomous exploration, navigation, and decision-making. In particular, the compact latent representations and rollout capabilities of ESWM make it a promising component for future embodied systems operating in unstructured physical environments.
Potential positive societal impacts include safer and more efficient navigation technologies, improved exploration strategies for robotics, and progress toward AI systems that reason about space in ways inspired by human cognition. At the same time, we recognize the risks associated with deploying autonomous agents in real-world settings, including the possibility of collecting sensitive data or influencing high-stakes decisions. These considerations underscore the importance of responsible development, careful evaluation in physical contexts, and adherence to ethical guidelines when integrating such models into embodied platforms.